# Preferences Evolve And So Should Your Bandits: Bandits with Evolving States for Online Platforms

### v2

@@ 12      Call Algorithm 2 for $\lambda=\hat{\lambda}_{1}$
After alternating between a random arm $i$ with arm $j$ for $(T/K)^{2/3}$ the state becomes (lemma E.3):
$q_{2(T/K)^{2/3}+1,(i,j)}=(1-\lambda)^{2(T/K)^{2/3}}q_{0,(i,j)}+\frac{1-(1-\lambda)^{2(T/K)^{2/3}}}{2-\lambda}b_{j}+(1-\lambda)\frac{1-(1-\lambda)^{2(T/K)^{2/3}}}{2-\lambda}b_{i}.$
And after that for another $(T/K)^{2/3}$ the algorithm samples reward from $\textsf{Bern}(r_{i}\cdot q_{(T/K)^{2/3}+2s+1})$, where
$q_{2(T/K)^{2/3}+2s+1,(i,j)}=(1-\lambda)^{2(T/K)^{2/3}+2s}q_{0,(i,j)}+\frac{1-(1-\lambda)^{2(T/K)^{2/3}+2s}}{2-\lambda}b_{j}+(1-\lambda)\frac{1-(1-\lambda)^{2(T/K)^{2/3}+2s}}{2-\lambda}b_{i}.$
for $s=\{0,...,(T/K)^{2/3}-1\}$. Thus, the expected reward $\mathbb{E}[R_{(i,j),s}]=r_{i}\cdot(1-\lambda)^{2(T/K)^{2/3}+2s}q_{0,(i,j)}$ $+\frac{1-(1-\lambda)^{2(T/K)^{2/3}+2s}}{2-\lambda}r_{i}b_{j}+(1-\lambda)\frac{1-(1-\lambda)^{2(T/K)^{2/3}+2s}}{2-\lambda}r_{i}b_{i}$. Thus,
$\mathbb{E}[\sum_{s=0}^{(T/K)^{2/3}-1}R_{(i,j),s}]$ $=\sum_{s=0}^{(T/K)^{2/3}-1}r_{i}\cdot(1-\lambda)^{2(T/K)^{2/3}+2s}q_{0,(i,j)}$
$+\frac{1-(1-\lambda)^{2(T/K)^{2/3}+2s}}{2-\lambda}r_{i}b_{j}+(1-\lambda)\frac{1-(1-\lambda)^{2(T/K)^{2/3}+2s}}{2-\lambda}r_{i}b_{i}$
$=(1-\lambda)^{2(T/K)^{2/3}}\left(r_{i}q_{0,(i,j)}+\frac{r_{i}b_{j}}{2-\lambda}+(1-\lambda)\frac{r_{i}b_{i}}{2-\lambda}\right)\frac{1-(1-\lambda)^{2(T/K)^{2/3}}}{1-(1-\lambda)^{2}}$
$+(T/K)^{2/3}\frac{r_{i}b_{j}+(1-\lambda)r_{i}b_{i}}{2-\lambda}.$
Thus, using Hoeffding’s inequality on $\hat{\mu}_{i,j}$ we get the result. ∎
If we take a union bound on all estimators we take that for all estimators $\hat{r}_{i}$ and $\hat{r}_{i,j}$ apply lemma E.5 and lemma E.6 with probability at least $2/T$ as $K<T$. In the continuation of the analysis we assume that applies.
@@
If $\lambda$ is in $[0,\tilde{\Theta}(1/T)]$ then the algorithm achieves regret:
@@
If $\lambda$ is in $[0,\tilde{\Theta}(1/T)]$ the states $q_{t}$ during the $K^{1/3}T^{2/3}$ rounds of building the estimators $\hat{\mu}_{i}$ and $\hat{\mu}_{i,j}$:
$q_{t}\geq(1-\lambda)^{K^{1/3}T^{2/3}}q_{0}\geq(1-1/T)^{K^{1/3}T^{2/3}}>1-\frac{1}{T^{1/3}}$
Thus, $|Y_{i}-Y_{(i,j)}|\leq r_{i}|q_{t}-q_{t}^{\prime}|$, with $t\in[0,K^{1/3}T^{2/3}]$ thus, $|q_{t}-q_{t}^{\prime}|\leq\frac{1}{T^{2/3}}$ and so $|Y_{i}-Y_{(i,j)}|\leq\frac{1}{T^{1/3}}$ for all arms $i,j$ in $[K]$. This means that $|\mu_{i}-\mu_{i,j}|\leq\frac{\sqrt{\log T}}{(T/K)^{1/3}}$ and after the $K\cdot(T/K)^{2/3}=K^{1/3}T^{2/3}$ rounds we run $EXP3.P$. From Section 4 we get :
$R_{\textsc{DES}}(T)=2K^{1/3}T^{2/3}+\mathcal{O}(\sqrt{KT\log T})+(1-(1-\lambda)^{T})\textsf{OPT}$
which ends the proof. ∎
@@
If $\max_{i,j\in[K]}\{|b_{i}-b_{j}|\}\leq\sqrt{\log T}/(T/K)^{1/3}$ then $\forall\lambda\in[0,1]$ Algorithm 8 after $K^{1/3}T^{2/3}$ calls $EXP3.P$ and achieves $R_{\textsc{DES}}(T)=\mathcal{O}\left(K^{1/3}T^{2/3}\right)$ regret.

### v4

@@ E.2.2 Analysis
At a high level, our proof proceeds as follows: (1) If the true $\lambda$ is near 00, then the state changes very little throughout the $T$ rounds; (2) the difference between $\hat{\mu}_{i}$ and $\hat{\mu}_{i,j}$ from Algorithm 8 is very small too; (3) Algorithm 8 calls EXP3.P.
Originally (i.e., in round $1$), the state is equal to $q_{1}=q_{0}=1$. Then, Algorithm 8 pulls all arms $i\in[K]$ for $K^{1/3}T^{2/3}$ rounds. If $\lambda$ is in $[0,\tilde{\Theta}(1/T)]$, the values of the states $q_{t}$ during the $K^{1/3}T^{2/3}$ rounds of building the estimators $\hat{\mu}_{i}$ and $\hat{\mu}_{i,j}$ will be close to each other. To see this, notice that: $1\geq(1-\lambda)^{K^{1/3}T^{2/3}}\geq(1-1/T)^{K^{1/3}T^{2/3}}>1-\frac{1}{T^{1/3}}$, which confirms that the state does not change significantly during $K^{1/3}T^{2/3}$ rounds. After a union bound over all arms $i\in[K]$ for both Lemmas E.4 and E.5, we get that with probability $1-2/T$:
$\hat{\mu}_{i}-\hat{\mu}_{i,j}\leq 2\frac{\sqrt{\log T}}{(T/K)^{1/3}}+\frac{1}{(T/K)^{2/3}}\mathbb{E}\left[\sum_{s=0}^{(T/K)^{2/3}-1}R_{s,i}\right]-\frac{1}{(T/K)^{2/3}}\mathbb{E}\left[\sum_{s=0}^{(T/K)^{2/3}-1}R_{s,i}^{\text{ALT}}\right]$
and also:
$\hat{\mu}_{i}-\hat{\mu}_{i,j}\geq-2\frac{\sqrt{\log T}}{(T/K)^{1/3}}+\frac{1}{(T/K)^{2/3}}\mathbb{E}\left[\sum_{s=0}^{(T/K)^{2/3}-1}R_{s,i}\right]-\frac{1}{(T/K)^{2/3}}\mathbb{E}\left[\sum_{s=0}^{(T/K)^{2/3}-1}R_{s,i}^{\text{ALT}}\right]$
Thus, to bound $|\hat{\mu}_{i}-\hat{\mu}_{i,j}|$ we want to bound the difference $\mathbb{E}\left[\sum_{s=0}^{(T/K)^{2/3}-1}R_{s,i}\right]-\mathbb{E}\left[\sum_{s=0}^{(T/K)^{2/3}-1}R_{s,i}^{\text{ALT}}\right]$:
$\mathbb{E}\left[\sum_{s=0}^{(T/K)^{2/3}-1}R_{s,i}\right]-\mathbb{E}\left[\sum_{s=0}^{(T/K)^{2/3}-1}R_{s,i}^{\text{ALT}}\right]$ $=\sum_{s=0}^{(T/K)^{2/3}-1}r_{i}\cdot\left((1-\lambda)^{(T/K)^{2/3}+s}\cdot(q_{0,i}-b_{i})+b_{i}\right)$
$\quad\quad-\sum_{s=0}^{(T/K)^{2/3}-1}r_{i}\cdot(1-\lambda)^{(T/K)^{2/3}+2s}\left(q_{0,i}^{\text{ALT}}-W_{i,j}(\lambda)\right)+r_{i}W_{i,j}(\lambda)$
$=\sum_{s=0}^{(T/K)^{2/3}-1}\underbrace{r_{i}\left((1-\lambda)^{(T/K)^{2/3}+s}q_{0,i}-(1-\lambda)^{(T/K)^{2/3}+2s}q_{0,i}^{ALT}\right)}_{A_{1}}$
$\quad\quad+\underbrace{r_{i}b_{i}\left(1-(1-\lambda)^{(T/K)^{2/3}+s}\right)}_{A_{2}}+\underbrace{r_{i}W_{i,j}\left(1-(1-\lambda)^{(T/K)^{2/3}+2s}\right)}_{A_{3}}$
Now remember that $(1-\lambda)^{K^{1/3}T^{2/3}}\geq 1-\frac{1}{(T/K)^{1/3}}$, where $K^{1/3}T^{2/3}$ are the rounds spent for building estimators $\hat{\mu}_{i}$s and $\hat{\mu}_{i,j}$s. Thus, the value of the state $q_{t}$ during these rounds takes values: $1-\frac{1}{(T/K)^{1/3}}\leq q_{t}\leq 1$, which implies that $-\Theta(1/(T/K)^{1/3})\leq A_{1}\leq\Theta(1/(T/K)^{1/3})$, $0\leq A_{2}\leq(1/(T/K)^{1/3})$ and $0\leq A_{3}\leq(1/(T/K)^{1/3})$. Hence,
$\frac{1}{(T/K)^{2/3}}\left|\mathbb{E}\left[\sum_{s=0}^{(T/K)^{2/3}-1}R_{s,i}\right]-\mathbb{E}\left[\sum_{s=0}^{(T/K)^{2/3}-1}R_{s,i}^{\text{ALT}}\right]\right|$ $\leq\Theta(1/(T/K)^{1/3}),$
and $|\hat{\mu}_{i}-\hat{\mu}_{i,j}|\leq 5\frac{\sqrt{\log T}}{(T/K)^{1/3}}$ with probability $1-2/T$. Conditioning on this event, Algorithm 8 will call EXP3.P (Algorithm 6). From Section 4, we get that the regret in this case is:
$R_{\textsc{DES}}(T)\leq\underbrace{2K^{1/3}T^{2/3}}_{\text{exploration samples}}+\underbrace{\mathcal{O}\left(\sqrt{KT\log T}\right)+\left(1-(1-\lambda)^{T}\right)\textsf{OPT}}_{\text{$R_{\textsc{DES}}$ of $\textsf{EXP3.P}$}}.$
Accounting for the regret picked up in the case where the event does not hold only adds an extra constant to the regret. This concludes our proof. ∎
Next, we bound the regret for $\lambda>\omega((K/T)^{2/3})$. To do this, we divide the analysis in two cases. In the first case (Lemma E.7), we have: $\max_{i,j\in[K]}\{r_{i}|b_{i}-b_{j}|\}\leq O(K^{1/3}T^{-1/3}\sqrt{\log T})$. Note that in this case, learning $\lambda$ (even approximately) is rather difficult. However, the difference between estimators $\hat{\mu}_{i}$ and $\hat{\mu}_{i,j}$ is rather small because by assumption the arms are not that different from each other. In the second case (Lemma E.8), we have $\max_{i,j\in[K]}\{r_{i}|b_{i}-b_{j}|\}>\omega(K^{1/3}T^{-1/3}\sqrt{\log T})$. For this case, Algorithm 8 is deployed in order to learn $\lambda$.
@@
If $\max_{i,j\in[K]}\{r_{i}|b_{i}-b_{j}|\}\leq O(\sqrt{\log T}/(T/K)^{1/3})$ and $\lambda>(K/T)^{2/3}$, then Algorithm 8 incurs regret $R_{\textsc{DES}}(T)=\widetilde{\mathcal{O}}\left(K^{1/3}T^{2/3}\right)$.
@@
Given that $\max_{i,j\in[K]}\{r_{i}|b_{i}-b_{j}|\}\leq O(\sqrt{\log T}/(T/K)^{1/3})$ and $\lambda>1/T^{2/3}$, we start by analyzing the regret of any algorithm (i.e., not necessarily EXP3.P) playing a sequence of actions $\{I_{t}\}_{t\in[T]}$:
$R_{\textsc{DES}}(T)$ $=\sum_{t\in[T]}r_{i^{{\pi^{\star}}}}q_{t}\left(H_{1:t-1}^{{\pi^{\star}}}\right)-\sum_{t\in[T]}r_{I_{t}}q_{t}\left(H_{1:t-1}^{\textsf{ALG}}\right)$
$=\sum_{t\in[T]}\left[(1-\lambda)^{t-1}q_{0}r_{i^{{\pi^{\star}}}}+r_{i^{{\pi^{\star}}}}\lambda\sum_{s=1}^{t}(1-\lambda)^{t-s}b_{i_{s}^{{\pi^{\star}}}}-(1-\lambda)^{t-1}q_{0}r_{I_{t}}+r_{I_{t}}\lambda\sum_{s=1}^{t}(1-\lambda)^{t-s}b_{I_{s}}\right]$
$=\sum_{t\in[T]}\left[(1-\lambda)^{t-1}q_{0}(r_{i^{{\pi^{\star}}}}-r_{I_{t}})+\sum_{s=1}^{t}(1-\lambda)^{t-s}(r_{i^{{\pi^{\star}}}}b_{i_{s}^{{\pi^{\star}}}}-r_{I_{t}}b_{I_{s}})\right]$
$\leq\sum_{t\in[T]}(1-\lambda)^{t-1}q_{0}(r_{i^{{\pi^{\star}}}}-r_{I_{t}})+O\left(\frac{\sqrt{\log T}}{(T/K)^{1/3}}\right),{}$ (34)
where the last inequality comes from the fact that $r_{i^{{\pi^{\star}}}}b_{i_{s}^{{\pi^{\star}}}}-r_{I_{t}}b_{I_{s}}$ is bounded by $O(\sqrt{\log T}/(T/K)^{1/3})$ for all $s\in[t]$ because $\max_{i,j\in[K]}\{r_{i}|b_{i}-b_{j}|\}\leq O(\sqrt{\log T}/(T/K)^{1/3})$.
For $\lambda>(K/T)^{2/3}$, $q_{0}=1$ and $|r_{i}-r_{j}|\leq 1,\forall i,j\in[K]$, the regret becomes:
$R_{\textsc{DES}}(T)$ $\leq\sum_{t\in[T]}(1-\lambda)^{t-1}q_{0}(r_{i^{{\pi^{\star}}}}-r_{I_{t}})+O\left(\frac{\sqrt{\log T}}{(T/K)^{1/3}}\right)$
$\leq\sum_{t\in[T]}(1-\lambda)^{t-1}+\widetilde{\mathcal{O}}(K^{1/3}T^{2/3})$
$\leq\frac{1-(1-\lambda)^{T}}{\lambda}+\widetilde{\mathcal{O}}(K^{1/3}T^{2/3})=T^{2/3}+\widetilde{\mathcal{O}}(K^{1/3}T^{2/3})$
$=\widetilde{\mathcal{O}}(K^{1/3}T^{2/3})$
@@
If $\max_{i,j\in[K]}\{r_{i}\cdot|b_{i}-b_{j}|\}>\omega(\sqrt{\log T}/(T/K)^{1/3})$ and $\lambda>\omega((K/T)^{2/3})$, then Algorithm 7 is called with $\tilde{N}$ such that: $\frac{N(\lambda)}{\log T}\leq\tilde{N}\leq 2N(\lambda)$ with probability at least $1-2(\log T+1)/T^{2}$.